# Finding Safety Neurons in Large Language Models

### v1

@@ 3.1 Investigation Setup
Models. To comprehensively investigate the safety neuron phenomenon in a more realistic setting, we use 3 different LLMs: Llama2-7b (Touvron et al., 2023), Mistral-7b-v0.1 (Jiang et al., 2023) and Gemma-7b (Team et al., 2024), denoted as Llama2, Mistral and Gemma for brevity, respectively. Details of these models can be found in appendix B.
@@
Alignment. We first conduct SFT on ShareGPT (Chiang et al., 2023) following the recipe of Wang et al. (2024). Then we perform safety alignment using DPO on the HH-RLHF-Harmless dataset (Bai et al., 2022a). We choose (IA)3 as our PEFT method and only apply it to the MLP layers. As (IA)3 functions by multiplying each activation by a re-scaling factor without changing their underlying parameters, this keeps the meanings of neurons unchanged, which is the basis of our method. The evaluation results of these models can be found in section E.2.
@@
Evaluation. We identify safety neurons on HH-RLHF-Harmless and evaluate the safety of LLMs on Beavertails (Ji et al., 2024). For metrics, we use the cost model beaver-7b-v1.0-cost from Dai et al. (2024). The cost model is a trained reward model that assigns a scalar score to each generation based on its safety (lower means safer). We will use cost score as our safety metric in the subsequent analysis due to its efficiency, widespread use, and alignment with human judgments (Liu et al., 2023; Duan et al., 2024; Kong et al., 2024). Further discussion on other evaluation metrics can be found in section E.1.
Figure 2: Cost score of patching three models (both Base and SFT version) with activations from DPO, while applied on top safety neurons and random neurons, evaluated on Beavertails.
3.2 Safety Neurons are Sparse and Effective
@@
Patching a large enough portion of neurons in activation patching can always restore the alignment performance. Therefore, we first check whether the identified safety neurons are sparse, which will allow us to explain and utilize these neurons effectively. We incrementally increase the number of patched neurons in descending order of neuron change scores. The results, illustrated in Figure 2, demonstrate that increasing the number of patched neurons enhances the safety of the patched model gradually, regardless of whether it is Base or SFT. Notably, after patching approximately $20,000$ neurons, SFT can recover over 90% of DPO’s safety performance, occasionally even exceeding the full DPO (Table 1). Consequently, we select the top $20,000$ neurons with the highest change scores (approximately 5.7%, 4.4%, and 2.9% for Llama2, Mistral, and Gemma, respectively) as investigated safety neurons in the subsequent experiments.
@@
To rule out the possibility that any arbitrary set of $20,000$ neurons from DPO can enhance model safety, we conduct experiments on randomly sampled neurons, ensuring that the number of neurons in each layer matches that of the safety neurons. The results, shown in Figure 2, indicate no causal effect of the randomly sampled neurons. We further conducted a t-test to compare the cost scores obtained from patching $20,000$ safety neurons versus random neurons. The p-values for all groups fall within the range from $1.15\times 10^{-6}$ to $1.67\times 10^{-18}$, indicating that the differences between random neurons and safety neurons are statistically significant. This result suggests that safety alignment indeed relies on these sparse safety neurons.
@@ 3.3 Safety Neurons Encode Transferable Mechanisms
Model BT ($\downarrow$) RT ($\downarrow$) HB ($\downarrow$) JL ($\downarrow$) PPL ($\downarrow$)
Llama2-7b Base $2.2$ $5.7$ $8.0$ $1.1$ $\mathbf{5.1}$
Base $-7.2$ $-5.5$ $-4.7$ $-8.3$ $5.6$
SFT $-2.4$ $-2.9$ $5.0$ $4.0$ $5.4$
SFT $\mathbf{-12.0}$ $\mathbf{-12.2}$ $-8.0$ $-7.6$ $5.4$
DPO $-11.8$ $-11.8$ $\mathbf{-11.0}$ $\mathbf{-10.5}$ $5.5$
Mistral-7b Base $-1.6$ $-4.8$ $-1.1$ $3.2$ $\mathbf{4.9}$
Base $-9.4$ $-10.1$ $\mathbf{-7.7}$ $\mathbf{-8.3}$ $5.1$
SFT $-7.6$ $-7.3$ $3.7$ $0.2$ $5.2$
SFT $-13.3$ $-12.6$ $-4.3$ $-6.0$ $5.3$
DPO $\mathbf{-13.5}$ $\mathbf{-13.4}$ $-6.1$ $-8.2$ $5.3$
Gemma-7b Base $1.1$ $0.4$ $7.8$ $1.1$ $\mathbf{6.6}$
Base $-8.1$ $-8.9$ $-1.2$ $-7.5$ $7.0$
SFT $-8.2$ $-9.8$ $1.0$ $-1.6$ $7.5$
SFT $-12.6$ $-12.7$ $-8.1$ $-8.5$ $7.6$
DPO $\mathbf{-13.6}$ $\mathbf{-14.1}$ $\mathbf{-11.9}$ $\mathbf{-10.6}$ $7.9$
Table 1: Cost scores on red-teaming datasets and perplexity on Wikitext-2. Abbr. BT = Beavertails, RT = RedTeam, HB = HarmBench, JL = JailBreakLLMs. denotes patching safety neurons’ activations from DPO.
We further investigate whether the effectiveness of safety neurons is transferrable by checking whether patching these neurons can enhance model safety on red-teaming benchmarks other than the trained dataset. To evaluate transferability, we select four benchmarks designed for red-teaming LLMs: Beavertails (Ji et al., 2024), RedTeam (Ganguli et al., 2022), HarmBench (Mazeika et al., 2024), and JailBreakLLMs (Shen et al., 2023). Following Ji et al. (2024), we also examine the influence on models’ general capability with perplexity on Wikitext-2 (Merity et al., 2016). The results, as shown in Table 1, indicate that the safety of the model improves significantly across all benchmarks after being patched with safety neuron activations. This demonstrates the transferability of safety neurons. Additionally, we observed that the perplexity of the patched model increased only marginally, and in most cases, the impact was less than that of DPO. This confirms safety neurons encode transferable mechanisms rather than shallow patterns depending on specific datasets.
Moreover, we investigate the related tokens of top safety neurons by projecting their corresponding value vectors into the vocabulary space (Geva et al., 2021), as shown in Table 2. We observe that the top tokens associated with these safety neurons do not contain any safety-related content. However, there are human-recognizable patterns among them, such as neurons promoting words related to food (third line in the table), conjunctions (fifth), and closing brackets (eighth). This differs from the toxic vectors identified by Lee et al. (2024), which suggests that reducing toxicity is done by avoiding the vectors related to toxic tokens. This difference may come from that our investigation range (comprehensive safety alignment) is larger than merely reducing toxicity. Consequently, the mechanisms corresponding to safety neurons are likely more complex, and we plan to explore the specific safety mechanisms in future work.
Vector Top Tokens
MLP.v3010106 ](#, ouc, iter, trat, ussen, tid, imos, ‖
MLP.v298343 </s>, Genomsnittlig, ]], ←, textt, <s>
@@
MLP.v267581 wa, ales, sin, MainActivity, oblig, raz
MLP.v299647 Food, Guard, Farm, Ali, Sex, Break, ob
@@
MLP.v284127 **, ».***, °, ”’, —-, /, !!, ]
MLP.v307219 Ż, Gemeinsame, HT, bez, Gor, category
[…]
3.4 Safety Neurons Emerge Stably
To further validate our findings, we explore whether safety neurons emerge stably in the alignment process, i.e., whether the randomness in the training process leads to identifying substantially different sets of safety neurons. We train five different SFT and DPO models using different random seeds and find that the overlap and Spearman’s rank correlation coefficients of the identified safety neurons exceed 95% across different models. Additionally, using these neurons to replicate the experiments in §3.2, we obtain the average cost scores and confidence intervals, as illustrated in Figure 3. These results indicate that the impact of training randomness on safety neurons is minimal.
Combining all these findings, we conclude that safety neurons are prevalent in the pre-trained base models, and safety alignment algorithms can leverage them to enhance LLMs’ safety, suggesting a possible mechanism of safety alignment. Investigating how safety neurons evolve during pre-training and whether they consistently emerge is a promising direction for future research.
Figure 3: Cost score of patching three models (both Base and SFT version) with activations from DPO on different number of safety neurons. The error bars are the 95% confidence interval over $5$ random trials.
@@ 4 Interpreting Alignment Tax
We first explore the relationship between safety neurons and other preference neurons, which are the neurons identified with our method for other preference-learning objectives. Specifically, we perform preference learning on 7 preference datasets: (1) Safety, including HH-Harmless (Harmless(Bai et al., 2022a) and RewardBench-Safety (Safety(Lambert et al., 2024); (2) Helpfulness, including HH-helpful (Helpful(Bai et al., 2022a) and Stanford Human Preferences (SHP(Ethayarajh et al., 2022); (3) Reasoning, including RewardBench-Reasoning (Reasoning(Lambert et al., 2024) and H4 Stack Exchange Preferences (H4SE(Lambert et al., 2023); (4) Information Extraction, including IEFeedback (Qi et al., 2024). Then, using the same method as for identifying safety neurons, we find the corresponding preference neurons, respectively, and calculate Spearman’s rank correlation coefficients between different preference neurons. The results are shown in Figure 4. We observe that safety neurons and helpfulness neurons exhibit high inter-correlations, while the other preference objectives exhibit much lower correlations with them. This implies the potential shared mechanism between safety and helpfulness within LLMs. The results of Mistral and Gemma can be found in section E.3.
Figure 4: Spearman’s rank correlation coefficients between preference neurons of Llama2 aligned on different preference-learning datasets.
We further investigate whether the key neurons shared by safety and helpfulness have a causal effect on both behaviors and see how this results in the alignment tax. We perform dynamic activation patching on the (around $15,000$) neurons shared between DPO on Harmless and DPO on Helpful and evaluate the influence on safety and helpfulness, which are evaluated on Beavertails using its cost model and reward model from Dai et al. (2024), respectively. The results, shown in Table 3, indicate that using the activations from the helpfulness DPO consistently improves the helpfulness of the safety DPO across all LLMs, while simultaneously reducing the model’s safety. The reverse direction yields similar results. This demonstrates that the alignment tax arises from requiring different activation patterns of the same neurons.
Patch Direction Safety ($\downarrow$) Helpfulness ($\uparrow$)
Llama2-7b
@@
Table 3: Performance changes of patching safety DPO models with activations of helpfulness DPO models on the neurons shared by safety and helpfulness, and vice versa. Safety and helpfulness are measured by cost and reward models, respectively. Green denotes performance decrease and Red denotes improvement.
@@ 5 Application: Safeguard for LLMs
We further explore the applications of our findings on safety neurons, presenting a preliminary use case: training a safeguard for LLMs based on safety neurons. The well-known Llama Guard (Inan et al., 2023) moderates LLM generations after detecting that harmful contents are generated, while we investigate whether the activations of safety neurons can predict harmful outputs before actual generation. This would enable us to reject harmful generation in advance, improving efficiency.
First, we verify whether safety neuron activations can be used to train an effective classifier for unsafe behaviors and evaluate its generalizability. We cache neuron activations at the last token of the prompt and create labels for these activations based on the cost scores of the corresponding generation (using a threshold of 0 to distinguish whether the generation is harmful or not) on the previously used 5 red-teaming benchmarks: HH-Harmless (Bai et al., 2022a), Beavertails (Ji et al., 2024), RedTeam (Ganguli et al., 2022), HarmBench (Mazeika et al., 2024), and JailBreakLLMs (Shen et al., 2023). A classifier trained on $1,500$ safety neuron activations from Beavertails achieves $79.4\%$ detection accuracy on average in the other benchmarks, indicating its potential for safeguarding LLMs.
We then test the safeguard performance. The safeguard is based on the detector trained on the HH-Harmless dataset, and we reject to output when harmful generations are detected. The average cost scores of the accepted responses are presented in Figure 5, which indicate that the safeguard obviously enhances the safety of unaligned models across all benchmarks. For models that have already undergone safety alignment, the safeguard can further improve their safety, thereby validating the application potential of this preliminary method. We provide additional experiments in section E.4.
Figure 5: Cost scores of different Mistral models on various red-teaming benchmarks.

### v2

@@ 4.1 Investigation Setup
Models. To comprehensively investigate the safety neuron phenomenon in a more realistic setting, we utilize four different pre-trained LLMs: Llama2-7b-hf (Touvron et al., 2023), Mistral-7b-v0.1 (Jiang et al., 2023), Gemma-7b (Team et al., 2024) and Qwen2.5-3B (Qwen et al., 2025), which we denote as Llama2, Mistral, Gemma and Qwen2.5 for brevity. Details of these models can be found in Appendix B.
@@
Alignment. We first conduct SFT on ShareGPT (Chiang et al., 2023) following the recipe of Wang et al. (2024). Then we perform safety alignment using DPO on the HH-RLHF-Harmless (Bai et al., 2022a). We select (IA)3(Liu et al., 2022) as our PEFT method and apply it exclusively to the MLP layers (details can be found in Appendix B.1). Since (IA)3 operates by multiplying each activation by a re-scaling factor without altering the underlying parameters, it preserves the functionality of the MLP neurons, which is fundamental to our approach as discussed before. The evaluation results of these models can be found in Appendix E.2.
@@
Evaluation. We compute change scores on HH-RLHF-Harmless and evaluate the causal effect on Beavertails (Ji et al., 2024). For metrics, we use the cost model beaver-7b-v1.0-cost from Dai et al. (2024). The cost model is a trained reward model that assigns a cost score to each prompt based on its safety (lower means safer). Then we compute the causal effect of neurons by Equation 4. We use cost score exclusively as our safety metric in the subsequent analysis due to its efficiency, widespread use, and alignment with human judgments (Liu et al., 2023; Duan et al., 2024; Kong et al., 2024). We also present the evaluation results using GPT-4 (Achiam et al., 2023) in Appendix E.1.
Baselines. Zhou et al. (2025) identified key attention heads whose ablation increases attack success rates (ASR), a finding that complements our work. However, due to differences in the focus of the two studies, we do not make a direct comparison. We use Pruning (Wei et al., 2024) and SN-Tune (Zhao et al., 2025) as baselines for comparison. Pruning employs methods such as the Wanda score (Sun et al., 2024) to locate parameters that have the greatest impact on ASR, while SN-Tune identifies safety neurons based on their contribution to residual stream on safety-related data. These ablation-based methods may be influenced by the existence of the "Hydra effect"(McGrath et al., 2023), which suggests that ablating certain components in LLMs may trigger compensatory behaviors in other components. The implementation details of these methods are described in Appendix B.3.
4.2 Safety Neurons are Sparse and Causally Effective
@@
Patching a large enough portion of neurons in activation patching can always restore the alignment performance. Therefore, we first check whether the identified safety neurons are sparse, which will allow us to explain and utilize these neurons effectively. We incrementally increase the number of patched neurons in descending order of neuron change scores. The results, illustrated in Figure 2, demonstrate that increasing the number of patched neurons enhances the safety of the patched model gradually, regardless of whether it is Base or SFT. Notably, after patching approximately $5\%$ of all the neurons, SFT can recover over $90\%$ of DPO’s safety performance, occasionally even exceeding the full DPO (Table 1).
@@
To rule out the possibility that patching any arbitrary set of neurons with activations DPO enhances model safety equally, we conduct experiments on randomly sampled neurons, ensuring that the number of neurons in each layer matches that of the safety neurons. The results, shown in Figure 2, indicate a significant causal effect gap between the randomly sampled neurons and safety neurons. We further conducted a t-test to compare the cost scores obtained from patching $5\%$ safety neurons versus random neurons. The p-values for all groups fall within the range from $1.15\times 10^{-6}$ to $1.67\times 10^{-18}$, indicating that the differences between random neurons and safety neurons are statistically significant. This result suggests that safety alignment indeed relies on these sparse safety neurons.
@@ 4.3 Safety Neurons Encode Transferable Mechanisms
Llama2 Base Mistral
Pruning +4+4 1-1 1-1 1-1 $\mathbf{0.00}$ 4-4 +3+3 +4+4 +4+4 $\mathbf{-0.01}$
SN-Tune 3-3 +1+1 5-5 +5+5 0.01-0.01 0 +12+12 4-4 +38+38 0.03-0.03
Ours $\mathbf{+56}$ $\mathbf{+65}$ $\mathbf{+63}$ $\mathbf{+78}$ 0.01-0.01 $\mathbf{+71}$ $\mathbf{+63}$ $\mathbf{+134}$ $\mathbf{+103}$ 0.04-0.04
SFT
Pruning +3+3 1-1 +4+4 1-1 $0.00$ 5-5 +7+7 4-4 6-6 +0.01+0.01
SN-Tune +13+13 +30+30 +11+11 +8+8 0.01-0.01 +10+10 +30+30 +21+21 +12+12 $0.00$
Ours $\mathbf{+101}$ $\mathbf{+101}$ $\mathbf{+76}$ $\mathbf{+73}$ $\mathbf{+0.01}$ $\mathbf{+90}$ $\mathbf{+80}$ $\mathbf{+74}$ $\mathbf{+75}$ $\mathbf{+0.01}$
Gemma Base Qwen2.5
Pruning +2+2 10-10 +6+6 +4+4 $\mathbf{0.00}$ 1-1 0 1-1 +5+5 $0.00$
SN-Tune 50-50 52-52 4-4 55-55 0.03-0.03 +30+30 +6+6 +3+3 +14+14 0.01-0.01
Ours $\mathbf{+78}$ $\mathbf{+68}$ $\mathbf{+64}$ $\mathbf{+70}$ 0.01-0.01 $\mathbf{+88}$ $\mathbf{+87}$ $\mathbf{+82}$ $\mathbf{+93}$ $\mathbf{+0.05}$
SFT
Pruning +9+9 +14+14 2-2 8-8 $\mathbf{0.00}$ 0 0 +4+4 0 $0.00$
SN-Tune 9-9 19-19 26-26 14-14 0.01-0.01 3-3 5-5 +4+4 3-3 $0.00$
Ours $\mathbf{+96}$ $\mathbf{+84}$ $\mathbf{+79}$ $\mathbf{+89}$ 0.02-0.02 $\mathbf{+83}$ $\mathbf{+81}$ $\mathbf{+68}$ $\mathbf{+83}$ $\mathbf{+0.01}$
Table 2: Value vectors of the top safety neurons from Llama2-7b, projected onto the vocabulary space. MLP.vllnn    denotes the down projection vector of the $n$-th neuron in layer $l$. We omitted some tokens for better visualization.
Vector Top Tokens
MLP.v285293 Sug, sugar, mouth, flesh
@@
MLP.v299647 Food, Guard, Farm, Break
@@
We further investigate whether the effectiveness of safety neurons is transferable by checking whether patching these neurons can enhance model safety on red-teaming benchmarks other than the trained datasets. To evaluate transferability, we select four benchmarks designed for red-teaming LLMs: Beavertails (Ji et al., 2024), RedTeam (Ganguli et al., 2022), HarmBench (Mazeika et al., 2024), and JailBreakLLMs (Shen et al., 2023). Additionally, we evaluate whether the enhancement of model safety comes at the expense of generation quality on various general benchmarks, including: Wikitext-2 (Merity et al., 2016), MMLU (Hendrycks et al., 2021), GSM8K (Cobbe et al., 2021), BBH (Suzgun et al., 2023), and TruthfulQA (Lin et al., 2022). We present the causal effects and average performance change to models without patching in Table 1, leaving the full original results in Appendix E.3. The results indicate that the safety of the model improves significantly across all benchmarks after being patched with safety neuron activations. This demonstrates the transferability of safety neurons. We also find the ablation-based methods are less causally effective in our verification setting. Additionally, we observed that the general capabilities of the patched model degenerated only marginally. This confirms that safety neurons encode transferable mechanisms rather than shallow patterns depending on specific datasets. The implementation details are described in B.2.
Moreover, we investigate the related tokens of top safety neurons by projecting their corresponding value vectors into the vocabulary space (Geva et al., 2021), as shown in Table 2 (full results are shown in Table 9). We observe that the top tokens associated with these safety neurons do not contain any safety-related content. However, there are human-recognizable patterns among them, such as neurons promoting words related to food, conjunctions, and closing brackets. This differs from the toxic vectors identified by Lee et al. (2024), which suggests that reducing toxicity is done by avoiding the vectors related to toxic tokens. This difference may come from our investigation range (comprehensive safety alignment) being larger than merely reducing toxicity. Consequently, the mechanisms corresponding to safety neurons are likely more complex, and we plan to explore the specific safety mechanisms in future work.
4.4 Safety Neurons Are Robust to Training Randomness
To further validate our findings, we explore whether safety neurons are robust in the alignment process, i.e., whether the randomness in the alignment training influences the identification of safety neurons. We train five different SFT and DPO models using different random seeds and find that the overlap and Spearman’s rank correlation coefficients of the identified safety neurons both exceed $0.95$ across different model families. Additionally, the error bars (Figure 2) obtained from repeating experiments in § 4.2 with these different models also indicate that the impact of training randomness on safety neurons is minimal.
Combining all these findings, we suggest that the safety neurons identified by our method are prevalent in the base models, and safety alignment algorithms exemplified by DPO (Rafailov et al., 2024) can moderate them to enhance LLMs’ safety, presenting a possible mechanism of safety alignment. Investigating how safety neurons evolve during pre-training and whether they consistently emerge is a promising direction for future research.
@@ 5 Interpreting Alignment Tax
[…]
We first explore the relationship between safety neurons and other preference neurons, which are the neurons identified with our framework for other preference-learning objectives. Specifically, we perform preference learning using DPO on 7 preference datasets categorized into 4 classes: (1) Safety, including HH-Harmless (Harmless(Bai et al., 2022a) and RewardBench-Safety (Safety(Lambert et al., 2024); (2) Helpfulness, including HH-helpful (Helpful(Bai et al., 2022a) and Stanford Human Preferences (SHP(Ethayarajh et al., 2022); (3) Reasoning, including RewardBench-Reasoning (Reasoning(Lambert et al., 2024) and H4 Stack Exchange Preferences (H4SE(Lambert et al., 2023); (4) Information Extraction, including IEFeedback (Qi et al., 2024). Then, using the same framework as for identifying safety neurons, we identify the top $5\%$ preference neurons respectively and calculate Spearman’s rank correlation coefficients between different preference neurons. The results of Llama2 are shown in Figure 3(a). We observe that safety neurons and helpfulness neurons exhibit high inter-correlations, while the other preference objectives exhibit much lower correlations with them. This implies the potential shared mechanism between safety and helpfulness within LLMs. The results of Mistral and Gemma can be found in Appendix E.4.
[…]
Patch Direction Safety Helpfulness
Llama2
@@
Qwen2.5
[…]
We further investigate whether the key neurons shared by safety and helpfulness have a causal effect on both behaviors and see how this results in the alignment tax. We perform dynamic activation patching between two DPOs trained on Harmless and Helpful with the preference neurons shared between models trained on Safety and SHP. We evaluate on Beavertails using its cost model and reward model from Dai et al. (2024), respectively. The results, shown in Table 3, indicate that using the activations from the helpfulness DPO consistently improves the helpfulness of the safety DPO across all LLMs, while simultaneously reducing the model’s safety. The reverse direction yields similar results. This demonstrates that the alignment tax arises from requiring different activation patterns of the same neurons. We further find that safety and helpfulness alignment indeed compete for control over these neurons during joint training (Appendix E.4). Besides, the causal effects of other preference neurons on safety and helpfulness (Figure 3(b)) are much lower, indicating different underlying mechanisms between safety/helpfulness and other capabilities.
@@ 6 Application: Safeguard for LLMs
We further explore the applications of our findings on safety neurons, presenting a preliminary use case: training a safeguard for LLMs based on safety neurons. The well-known Llama Guard (Inan et al., 2023) moderates LLM generations after detecting that harmful contents are generated, while we investigate whether the activations of safety neurons can predict harmful outputs before actual generation. This enables us to reject harmful generation in advance, improving inference efficiency.
Figure 4: Cost scores (linear transformed for better visualization) of four models with safeguard on red-teaming benchmarks.
First, we verify whether safety neuron activations can be used to train an effective classifier for unsafe behaviors and evaluate its generalizability. We cache neuron activations from SFT at the last token of the prompt and create labels for these activations based on the cost scores of the corresponding generation  [^1] We use a threshold of 0 to distinguish whether the generation is harmful or not. on the previously used 5 red-teaming benchmarks: HH-Harmless (Bai et al., 2022a), Beavertails (Ji et al., 2024), RedTeam (Ganguli et al., 2022), HarmBench (Mazeika et al., 2024), and JailBreakLLMs (Shen et al., 2023). A comprehensive cross-validation demonstrates the classifier, trained on $1500$ safety neuron activations, achieves $76.2\%$ accuracy on average, indicating its potential for safeguarding LLMs. More detailed results are in Appendix E.5.
We can use the trained classifier to predict whether the LLM will produce harmful content before generating the first token. Then, we can either halt generation and output a predefined response or continue generating with a refusal prefix (e.g., ‘sorry’). We apply the safeguard trained on SFT activations from HH-Harmless to both SFT and DPO, with a simple evaluation protocol: we compute the average cost scores on accepted responses as a proxy for safeguarding results. The results, presented in Figure 4, indicate that the safeguard significantly enhances the safety of unaligned models across all benchmarks. For models that have already undergone safety alignment, the safeguard can further improve safety. Besides, the overhead of the classifier is less than $0.001$ seconds, which is negligible compared to generation, validating the potential value of this preliminary method.